OpenStudy (jamie.):

How do you do X-4y=4

alones (alones):

D oyou need to solve it or graph? So i solved for x x−4y=4 add 4y to both sides. x−4y+4y=4+4y x=4y+4 should be your asnwer or if you need this ot be graphed then https://www.desmos.com/calculator

OpenStudy (jamie.):

Thank you

alones (alones):

Yo welcome